# Investigations of MWISP Clumps: 13CO Clump Source Catalogs and Physical Properties

### v1

@@ 1 Introduction
Molecular clouds in the interstellar medium exhibit rich and complex structures across multiple spatial scales, serving as critical components in galactic evolution. Within this hierarchical organization, molecular clumps represent intermediate-scale condensations that mark a crucial transition from diffuse gas to star formation (Bergin & Tafalla, 2007; Hacar et al., 2013; Motte et al., 2018; Ballesteros-Paredes et al., 2020). These clumps form through complex interactions between gravity, turbulence, magnetic fields, and stellar feedback, often occurring within the filamentary structures of larger molecular cloud complexes (e.g. Hennebelle & Falgarone, 2012; André et al., 2014; Bally, 2016; Hennebelle & Inutsuka, 2019; Pineda et al., 2023). The diverse formation pathways—whether through gravitational instability, turbulent compression, or filamentary fragmentation—concentrate gas into high-density regions capable of gravitational collapse. Characterizing clump properties thus provides direct insights into both the initial conditions of star formation and the critical balance between collapsing forces and supportive mechanisms like turbulent pressure or magnetic tension (e.g. Krumholz et al., 2005; Vázquez-Semadeni et al., 2019; Hennebelle & Inutsuka, 2019), ultimately revealing the physical processes that govern molecular cloud evolution across galactic environments (e.g. Krumholz, 2014; Motte et al., 2018).
While decades of surveys have mapped molecular cloud distributions, comprehensive studies of clump properties across the Galaxy remain limited(e.g. Sanders et al., 1986; Dame et al., 1987, 2001; Heyer & Dame, 2015). The Boston University-Five College Radio Astronomy Observatory Galactic Ring Survey (GRS; Jackson et al. 2006) contributed important 13CO observations but with limited spatial coverage, while dust continuum surveys like ATLASGAL (Schuller et al., 2009) and Hi-GAL (Molinari et al., 2010) offered comprehensive coverage of high-density regions but lacked velocity information necessary for distance determination and dynamical analysis. More recent efforts, including ThrUMMS (Barnes et al., 2015), CHIMPS (Rigby et al., 2016), FUGIN (Umemoto et al., 2017), SEDIGISM (Schuller et al., 2017), FQS (Benedettini et al., 2020), OGHReS (Urquhart et al., 2024), and PAMS (Rigby et al., 2025) have advanced our understanding of molecular cloud structures but still face limitations in spatial coverage, sensitivity, or ability to trace gas across different density regimes.
@@
The Milky Way Imaging Scroll Painting (MWISP, Su et al. 2019) project offers a transformative approach to addressing these limitations through its comprehensive, unbiased survey of 12CO/13CO/C18O ($J=1\text{--}0$) emission across the Galactic plane. Spanning an wide-ranging region in its Phase I operations, MWISP represents one of the most spatially complete molecular line surveys of the Milky Way to date. The simultaneous acquisition of multiple CO isotopologues enables a multi-tracer perspective that captures molecular gas across diverse density regimes and environmental conditions. This coordinated multi-isotopologue strategy allows for precise determination of critical physical parameters, such as excitation temperatures, optical depths, and column densities, through local thermodynamic equilibrium (LTE) analysis (Rigby et al., 2019; Dong et al., 2023).
@@
The paper is organized as follows: Section 2 introduces the MWISP survey and details the observational data utilized in this study. Section 3 presents the molecular clump detection methodology, including the FacetClumps algorithm and the equivalent global detection scheme. Section 4 details the hierarchical approach to distance determination and resolution of kinematic distance ambiguities. Section 5 analyzes the physical properties and scaling relations of molecular clumps, comparing our results with other major surveys and examining differences in maser-associated clump populations. Finally, Section 6 presents the conclusions and outlines future research directions enabled by these clump catalogs.
@@ 2.1 The MWISP Survey and 13CO Emission
Figure 1: Schematic diagram illustrating the equivalent global detection scheme. (a) The SRs extracted from the original data (see Figure 1 of Jiang et al. 2025) by FacetClumps. Curves of different colors represent the boundaries of different SRs. (b) The reconstructed SR data cube, corresponding to the red contours in (a). The asterisks denote the centers of the clumps, and the direction of the red straight lines indicates the orientation of the clumps at the untouched edges. Contours in different colors represent the boundaries of the clumps. To avoid confusion, only a subset of the clump boundaries is shown. (c) The reconstructed SR data cube, corresponding to the blue contours in (a). (d) Detection results storage data cube. Transfer the detection masks from (b) and (c) to a storage data cube of the same size as the original data. The inset view shows two clumps in details, and the black circle in the bottom-left corner marks the spatial resolution of MWISP.
@@ 3.1 The Clump Detection Algorithm
Molecular clumps, as dense conglomerations of gas and dust within molecular clouds, are defined by a central region with conspicuously higher density than their surrounding environment (Bergin & Tafalla, 2007; Jiang et al., 2023). This characteristic makes them critical potential sites for star formation. Morphologically, these clumps exhibit irregular geometries (Luo et al., 2024a; Long et al., 2024), and their average intensity spectrum along any direction exhibits a Gaussian-like profile.
Accurate identification of such clumps is essential to understanding star formation processes. The FacetClumps algorithm (Jiang et al., 2023) has been demonstrated to exhibit robust performance in localization accuracy, detection completeness, and regional segmentation precision (Berry, 2015; Luo et al., 2022; Jiang et al., 2022). Initially, FacetClumps extracts signal regions (SRs) using morphological methods. By combining the facet model with the extremum determination theorem of multivariate functions, it autonomously locates clump centers within the preprocessed SRs. The algorithm then applies a connectivity-based minimum distance clustering method to merge locally gradient-segmented regions, effectively delineating the boundaries of individual clumps. Given its technical advantages and demonstrated reliability, this study utilizes FacetClumps for clump detection. Detailed parameter configurations are presented in Table 2 of Appendix A.
@@ 3.2 The Equivalent Global Detection Scheme
Our study seeks to achieve direct global detection across entire 2310 deg2 observational datasets. However, such global detection requires prohibitively high computational resources that become increasingly challenging with growing data volumes. Traditional strategies using subdivided data cubes – even with overlapping regions partitioning – suffer from two critical limitations: (1) mask stitching ambiguities at subcube boundaries, and (2) selection ambiguities for clumps in overlapping areas. These issues become particularly severe for extended structures that span multiple subcubes, where discrepancies between subcube-level and global-level detection are inevitable. In such cases, even a single misidentified clump can trigger cascading mask stitching errors.
To overcome these challenges, we propose an equivalent global detection scheme outlined in Figure 1. The original illustration data and direct detection results have been presented in Figures 1 and 2 of Jiang et al. (2025). The implementation consists of three stages: (1) Initial SR extraction using the first detection phase of FacetClumps (Figure 1a), ensuring the extraction of the same objects across different-scale detections; (2) Independent clump detection for each SR (Figure 1b-c), with simultaneous mapping of generated masks to a global storage array matching the size of the original data (Figure 1d); (3) Final catalog calculation using original data and integrated global masks. The magnified inset demonstrates the characteristic morphologies of clumps and confirms they exceed the telescope’s beam-scale threshold required for physically meaningful source extraction.
@@
[…]
3.3 Clumps Information and Detection Catalog
The MWISP Phase I, spanning a decade of observations, maintains generally uniform sensitivity across its datasets, yet localized variations emerged due to temporal changes in observational parameters (e.g., atmospheric transparency fluctuations and receiver configuration upgrades). To mitigate these variations, we perform spectral-axis-normalized noise calibration through pixel-wise division of the raw data cube by its corresponding noise map, generating a unified signal-to-noise ratio (S/R) cube (where the noise RMS = 1). This S/R-based processing approach optimally preserves high-S/R sources in low-background regions while suppressing false detections in high-noise areas (e.g. Eden et al., 2017; Rigby et al., 2019; Rani et al., 2023).
We implement the equivalent global detection scheme in conjunction with FacetClumps on the S/N cube, whereby the S/N cube serves exclusively as the basis for mask extraction. Clump property calculations utilize the original brightness-temperature data with the corresponding global mask. We have extracted 71,661 molecular clumps within approximately 6 hours of computational time. Representative clump images are displayed in Figure 16 of Appendix E, and the detection catalog is detailed in Table 3 of Appendix B.
@@ 4 Distance Determination
Accurate distance determination to clumps forms the cornerstone for deriving their physical properties. While distances to a number of nearby large-scale molecular clouds have been established using reddened background stars, these measurements continue to undergo updates and refinements (Yan et al., 2019, 2020, 2021; Mei et al., 2024; Sun et al., 2024a; Zhang et al., 2024; Zhuang et al., 2024). To overcome the challenge of distance determination for broad sky-survey samples such as MWISP, we introduce a hierarchical distance decision method that prioritizes maser parallax distances when available and incorporating kinematic distance measurements (Urquhart et al., 2018), ensuring maximum completeness in distance determination across the entire sample.
@@
4.1 Resolving KDA of SRs
The KDA problem in the FGQ SR is resolved using the process outlined in Figure 4. This process consists of seven main subprocesses, with the reliability of each step gradually decreasing. Before making determinations, the kinematic near, far, and tangent point distances are calculated for each SR and clump using the Galactic rotation curve model ($R_{0}=8.15$ kpc and $\theta_{0}=240$ km s-1, Wenger et al. 2018; Reid et al. 2019).
@@ 4.1 Resolving KDA of SRs
  • Height Constraint (4): When a far distance solution results in a height $Z>120~\mathrm{pc}$ (the galactic disk height threshold of clumps, Urquhart et al., 2011a, 2018; Su et al., 2021), the SR is classified as near distance.
@@
    We have compiled HI surveys overlapping with the MWISP survey in the FGQ, prioritizing data sources based on resolution: the Galactic Arecibo L-Band Feed Array Hi survey (GALFA-Hi, Peek et al., 2011), the VLA Galactic Plane Survey (VGPS, Stil et al., 2006), the Southern Galactic Plane Survey (SGPS, McClure-Griffiths et al., 2005), the Canadian Galactic Plane Survey (CGPS, Taylor et al., 2003), and HI4PI (HI4PI Collaboration et al., 2016). Figure 5 illustrates the distribution of SRs determined by HISA from these various projects.
    To systematically and objectively evaluate HISA features across such a broad sky area, we developed an automated classification method illustrated in Figure 6. When self-absorption features are detected near the SR, the distance is classified as near. Conversely, the presence of HI emission peaks suggests a far classification. If neither feature is apparent, the distance remains undetermined.
@@
    Figure 6: Example of automatic HISA identification for the source G18.107+0.409+22.913 (filament Fil 7 from Jiang et al., 2025). Left Panel: The black curve shows the signal-point HI spectrum of the target source from the VGPS (Stil et al., 2006) survey, with the orange vertical line indicating the velocity of the source. Right Panel: This is a zoomed-in view of the left panel. Red points mark the local maxima of the HI profile, while blue points mark minima. The velocity range between adjacent extrema is divided into three segments: the red segment (near maximum) corresponds to the far kinematic distance solution, the blue segment (near minimum) corresponds to the near solution, and the green intermediate segment denotes ambiguous cases. The target’s systemic velocity (orange line) falls within the blue segment. The velocity difference between the target and the nearest minimum (1.182 km s-1) is smaller than the tolerance threshold (3 km s-1), leading to its classification as the near kinematic distance at this spectrum.
    The automated approach first identifies all local maximum and minimum in the HI spectral line and divides the velocity range of local extrema around the target velocity into three equal segments. The distance is classified as near when the target velocity falls within a local minimum region with a velocity difference smaller than the tolerance (typically 3 km s-1, Mège et al., 2021). When the target velocity occurs within a local maximum region with the same velocity tolerance, the classification is far. Target velocities falling in the intermediate range yield an unknown classification.
@@
  • Matching IRDCs (6): The kinematic distance assignment for Infrared Dark Clouds (IRDCs) relies on their mid-infrared extinction features against Galactic background emission (Simon et al., 2006). Since detecting extinction requires mid-infrared radiation from the inner Galactic disk, this method tends to favor near-distance solutions (Giannetti et al., 2015; Urquhart et al., 2018; Duarte-Cabral et al., 2021; Mège et al., 2021), with an associated uncertainty of approximately 10% (Giannetti et al., 2015).
    If an IRDC lies within an SR, and the SR at the center of the IRDC exhibits the most velocity channels along the line of sight, the SR is classified as near. Due to line-of-sight blending, an IRDC may correspond to multiple SRs at different velocities; however, IRDCs generally correspond to the strongest velocity component (e.g. Kainulainen & Tan, 2013). This approach helps minimize the interference of weaker components by the stronger ones, thereby enhancing the accuracy of the classification.
  • Parameter Relations Method (7): Distance determination based on the empirical $\sigma_{v}$-$R$ relation (Larson, 1981) resolves near and far kinematic ambiguities by requiring consistency with established scaling laws, a methodology widely implemented in inner Galactic molecular cloud studies (Dame & Thaddeus, 1985; Solomon et al., 1987; Duarte-Cabral et al., 2021). However, as demonstrated in Section 5.3.1 and corroborated by previous investigations (e.g., Heyer et al., 2009; Heyer & Dame, 2015; Miville-Deschênes et al., 2017; Feng et al., 2024), the composite parameter relation $\sigma_{v}$$\Sigma R$ exhibits reduced scatter and demonstrates enhanced robustness across diverse dynamical environments within the Milky Way. Accordingly, we adopt this scaling relation as the premise for developing near or far distance discriminants (Miville-Deschênes et al., 2017).
    Building upon obtained SR classifications, we propagate distance solutions to individual clumps through the process detailed in Section 4.2, successfully determining distances for over 96% of clumps in the FGQ. We construct a $\sigma_{v}$-$R\Sigma$ distribution plane using samples with known distances as described in Section 5.3.1. For each clump with undetermined classification, we calculate the Mahalanobis distance (Virtanen et al., 2020) between its derived ($\sigma_{v}$, $R\Sigma$) parameters and the multivariate distributions corresponding to both near and far kinematic solutions. This statistical metric, superior to Euclidean alternatives due to its covariance-matrix normalization, effectively accounts for intrinsic parameter correlations in the $\sigma_{v}$-$R\Sigma$ plane and accommodates varying measurement uncertainties. We then adopt the kinematic solution that minimizes this metric.
@@
Figure 8: Face-on number density distribution map of clumps in the Galactic plane. The coordinate system centers on the Galactic Center (red star) with the Sun positioned at $(0,8.15)$ kpc (yellow circle). The red dash-dot circle represents the 3 kpc ring around the Galactic Center, while the blue dash-dot circle represents the Solar Circle. The number density distribution is calculated using KDE, where the color scale represents the density of clumps, with higher density regions depicted in greater estimated values. Gray dashed lines show Galactic longitude lines (10° intervals). Axes are centered at the Galactic Center, with negative Y-axis toward $l=0^{\circ}$ and positive X-axis toward $l=90^{\circ}$. The colored dashed lines depict the spiral arms, with colors matching those in the bottom panel of Figure 2, based on the model by Reid et al. (2019).

### v2

@@ 1 Introduction
Molecular clouds in the interstellar medium exhibit rich and complex structures across multiple spatial scales, serving as critical components in galactic evolution. Within this hierarchical organization, molecular clumps represent intermediate-scale condensations that mark a crucial transition from diffuse gas to star formation (Bergin & Tafalla, 2007; Hacar et al., 2013; Motte et al., 2018; Ballesteros-Paredes et al., 2020). These clumps form through complex interactions between gravity, turbulence, magnetic fields, and stellar feedback, often occurring within the filamentary structures of larger molecular cloud complexes (e.g., Hennebelle & Falgarone, 2012; André et al., 2014; Bally, 2016; Hennebelle & Inutsuka, 2019; Pineda et al., 2023). The diverse formation pathways—whether through gravitational instability, turbulent compression, or filamentary fragmentation—concentrate gas into high-density regions capable of gravitational collapse. Characterizing clump properties thus provides direct insights into both the initial conditions of star formation and the critical balance between collapsing forces and supportive mechanisms like turbulent pressure or magnetic tension (e.g. Krumholz et al., 2005; Vázquez-Semadeni et al., 2019; Hennebelle & Inutsuka, 2019), ultimately revealing the physical processes that govern molecular cloud evolution across galactic environments (e.g. Krumholz, 2014; Motte et al., 2018).
While decades of surveys have mapped molecular cloud distributions, comprehensive studies of clump properties across the Galaxy remain limited (e.g., Sanders et al., 1986; Dame et al., 1987, 2001; Heyer & Dame, 2015). The Boston University-Five College Radio Astronomy Observatory Galactic Ring Survey (Jackson et al. 2006) contributed important 13CO observations but with limited spatial coverage, while dust continuum surveys like ATLASGAL (Schuller et al., 2009) and Hi-GAL (Molinari et al., 2010) offered comprehensive coverage of high-density regions but lacked velocity information necessary for distance determination and dynamical analysis. More recent efforts, including ThrUMMS (Barnes et al., 2015), CHIMPS (Rigby et al., 2016), FUGIN (Umemoto et al., 2017), SEDIGISM (Schuller et al., 2017), FQS (Benedettini et al., 2020), OGHReS (Urquhart et al., 2024), and PAMS (Rigby et al., 2025) have advanced our understanding of molecular cloud structures but still faced limitations in spatial coverage, sensitivity, or ability to trace gas across different density regimes.
@@
The Milky Way Imaging Scroll Painting (MWISP; Su et al. 2019) project offers a transformative approach to addressing these limitations through its comprehensive, unbiased survey of 12CO/13CO/C18O ($J=1\text{--}0$) emission across the Galactic plane. Spanning an wide-ranging region in its Phase I operations, MWISP represents one of the most spatially complete molecular line surveys of the Milky Way to date. The simultaneous acquisition of multiple CO isotopologues enables a multi-tracer perspective that captures molecular gas across diverse density regimes and environmental conditions. This coordinated multi-isotopologue strategy allows for precise determination of critical physical parameters, such as excitation temperatures, optical depths, and column densities, through local thermodynamic equilibrium (LTE) analysis (Rigby et al., 2019; Dong et al., 2023).
@@
The paper is organized as follows: Section 2 introduces the MWISP survey and details the observational data utilized in this study. Section 3 presents the molecular clump detection methodology, including the FacetClumps algorithm and the equivalent global detection scheme. Section 4 details the hierarchical approach to distance determination and resolution of KDAs. Section 5 analyzes the physical properties and scaling relations of molecular clumps, comparing our results with other major surveys and examining differences in maser-associated clump populations. Finally, Section 6 presents the conclusions and outlines future research directions enabled by these clump catalogs.
@@ 2.1 The MWISP Survey and 13CO Emission
Figure 1: Schematic diagram illustrating the equivalent global detection scheme. (a) The SRs extracted from the original data (see Figure 1 of Jiang et al. 2025) by FacetClumps. Curves of different colors represent the boundaries of different SRs. (b) The reconstructed SR data cube, corresponding to the red contours in (a). The asterisks denote the centers of the clumps, and the direction of the red straight lines indicates the orientation of the clumps at the untouched edges. Contours in different colors represent the boundaries of the clumps. To avoid confusion, only a subset of the clump boundaries is shown. (c) The reconstructed SR data cube, corresponding to the blue contours in (a). (d) Detection results storage data cube. Transfer the detection masks from (b) and (c) to a storage data cube of the same size as the original data. The inset view shows two clumps in detail, and the black circle in the bottom-left corner marks the spatial resolution of MWISP.
@@ 3.1 The Clump Detection Algorithm
Molecular clumps, as dense conglomerations of gas and dust within molecular clouds, are defined by a central region with conspicuously higher density than the surrounding environment (Bergin & Tafalla, 2007; Jiang et al., 2023). This characteristic makes them critical potential sites for star formation. Morphologically, these clumps exhibit irregular geometries (Long et al., 2024; Luo et al., 2024a), and their average intensity spectrum along any direction exhibits a Gaussian-like profile.
Accurate identification of such clumps is essential to understanding star formation processes. The FacetClumps algorithm (Jiang et al., 2023) has been demonstrated to exhibit robust performance in localization accuracy, detection completeness, and regional segmentation precision (Berry, 2015; Jiang et al., 2022; Luo et al., 2022). Initially, FacetClumps extracts signal regions (SRs) using morphological methods. By combining the facet model with the extremum determination theorem of multivariate functions, it autonomously locates clump centers within the preprocessed SRs. The algorithm then applies a connectivity-based minimum distance clustering method to merge locally gradient-segmented regions, effectively delineating the boundaries of individual clumps. Given its technical advantages and demonstrated reliability, this study utilizes FacetClumps for clump detection. Detailed parameter configurations are presented in Table A.1 of Appendix A.
@@ 3.2 The Equivalent Global Detection Scheme
Our study seeks to achieve direct global detection across entire 2310 deg2 observational datasets. However, such global detection requires prohibitively high computational resources that become increasingly challenging with growing data volumes. Traditional strategies using subdivided data cubes – even with overlapping region partitioning – suffer from two critical limitations: (1) mask stitching ambiguities at subcube boundaries, and (2) selection ambiguities for clumps in overlapping areas. These issues become particularly severe for extended structures that span multiple subcubes, where discrepancies between subcube-level and global-level detection are inevitable. In such cases, even a single misidentified clump can trigger cascading mask stitching errors.
To overcome these challenges, we propose an equivalent global detection scheme outlined in Figure 1. The original illustration data and direct detection results have been presented in Figures 1 and 2 of Jiang et al. (2025). The implementation consists of three stages: (1) initial SR extraction using the first detection phase of FacetClumps (Figure 1a), ensuring the extraction of the same objects across different-scale detections; (2) independent clump detection for each SR (Figure 1b–c), with simultaneous mapping of generated masks to a global storage array matching the size of the original data (Figure 1d); (3) final catalog calculation using original data and integrated global masks. The magnified inset demonstrates the characteristic morphologies of clumps and confirms they exceed the telescope’s beam-scale threshold required for physically meaningful source extraction.
@@
Figure 3: Distribution of clumps across the Galactic longitude–latitude plane. Box plots illustrate the statistical distribution of Galactic latitude within discrete longitude intervals. Each box displays the interquartile range (IQR), with upper and lower boundaries marking the third and first quartiles, respectively, and the horizontal line indicating the median. The colored dashed-dotted lines represent median values for specific populations: red for positive Galactic latitudes (GLat >> 0), blue for negative latitudes (GLat << 0), and black for the overall sample. Whiskers extend to the most extreme data points within 1.5 $\times$ IQR, revealing the dispersion across different Galactic longitudes. The orange line (right axis) quantifies the source count within each longitude bin.
3.3 Clump Information and Detection Catalog
The MWISP Phase I, spanning a decade of observations, maintains generally uniform sensitivity across its datasets, yet localized variations emerged due to temporal changes in observational parameters (e.g., atmospheric transparency fluctuations and receiver configuration upgrades). To mitigate these variations, we perform spectral-axis-normalized noise calibration through pixel-wise division of the raw data cube by its corresponding noise map, generating a unified signal-to-noise ratio (S/R) cube (where the noise RMS = 1). This S/R-based processing approach optimally preserves high-S/R sources in low-background regions while suppressing false detections in high-noise areas (e.g., Eden et al., 2017; Rigby et al., 2019; Rani et al., 2023).
We implement the equivalent global detection scheme in conjunction with FacetClumps on the S/N cube, where the S/N cube serves exclusively as the basis for mask extraction. Clump property calculations utilize the original brightness temperature data with the corresponding global mask. We have extracted 71,661 molecular clumps within approximately 6 hours of computational time. Representative clump images are displayed in Figure E.1 of Appendix E, and the detection catalog is detailed in Table B.1 of Appendix B.
@@ 4 Distance Determination
Accurate distance determination to clumps forms the cornerstone for deriving their physical properties. While distances to a number of nearby large-scale molecular clouds have been established using reddened background stars, these measurements continue to undergo updates and refinements (Yan et al., 2019, 2020, 2021; Mei et al., 2024; Sun et al., 2024a; Zhang et al., 2024; Zhuang et al., 2024). To overcome the challenge of distance determination for broad sky-survey samples such as MWISP, we introduce a hierarchical distance decision method that prioritizes maser parallax distances when available and incorporates kinematic distance measurements (Urquhart et al., 2018), ensuring maximum completeness in distance determination across the entire sample.
@@
4.1 Resolving the KDA of SRs
The KDA problem in the FGQ SR is resolved using the process outlined in Figure 4. This process consists of seven main subprocesses, with the reliability of each step gradually decreasing. Before making determinations, the kinematic near, far, and tangent point distances are calculated for each SR and clump using the Galactic rotation curve model ($R_{0}=8.15$ kpc and $\theta_{0}=240$ km s-1; Wenger et al. 2018; Reid et al. 2019).
@@ 4.1 Resolving the KDA of SRs
  • Height constraint (4): When a far-distance solution results in a height $Z>120~\mathrm{pc}$ (the galactic disk height threshold of clumps; Urquhart et al., 2011a, 2018; Su et al., 2021), the SR is classified as near distance.
@@
    We have compiled HI surveys overlapping with the MWISP survey in the FGQ, prioritizing data sources based on resolution: the Galactic Arecibo $L$-Band Feed Array Hi survey (GALFA-Hi; Peek et al., 2011), the VLA Galactic Plane Survey (VGPS; Stil et al., 2006), the Southern Galactic Plane Survey (SGPS; McClure-Griffiths et al., 2005), the Canadian Galactic Plane Survey (CGPS; Taylor et al., 2003), and HI4PI (HI4PI Collaboration et al., 2016). Figure 5 illustrates the distribution of SRs determined by HISA from these various projects.
    To systematically and objectively evaluate HISA features across such a broad sky area, we develope an automated classification method illustrated in Figure 6. When self-absorption features are detected near the SR, the distance is classified as near. Conversely, the presence of HI emission peaks suggests a far classification. If neither feature is apparent, the distance remains undetermined.
@@
    Figure 6: Example of automatic HISA identification for the source G18.107+0.409+22.913 (filament Fil 7 from Jiang et al., 2025). Left panel: the black curve shows the signal-point HI spectrum of the target source from the VGPS (Stil et al., 2006) survey, with the orange vertical line indicating the velocity of the source. Right panel: this is a zoomed-in view of the left panel. Red points mark the local maxima of the HI profile, while blue points mark minima. The velocity range between adjacent extrema is divided into three segments: the red segment (near maximum) corresponds to the far kinematic distance solution, the blue segment (near minimum) corresponds to the near solution, and the green intermediate segment denotes ambiguous cases. The target’s systemic velocity (orange line) falls within the blue segment. The velocity difference between the target and the nearest minimum (1.182 km s-1) is smaller than the tolerance threshold (3 km s-1), leading to its classification as near kinematic distance at this spectrum.
    The automated approach first identifies all local maxima and minima in the HI spectral line and divides the velocity range of local extrema around the target velocity into three equal segments. The distance is classified as near when the target velocity falls within a local minimum region with a velocity difference smaller than the tolerance (typically 3 km s-1; Mège et al., 2021). When the target velocity occurs within a local maximum region with the same velocity tolerance, the classification is far. Target velocities falling within the intermediate range yield an unknown classification.
@@
  • Matching Infrared Dark Clouds (IRDCs) (6): The kinematic distance assignment for IRDCs relies on their mid-infrared extinction features against Galactic background emission (Simon et al., 2006). Since detecting extinction requires mid-infrared radiation from the inner Galactic disk, this method tends to favor near-distance solutions (Giannetti et al., 2015; Urquhart et al., 2018; Duarte-Cabral et al., 2021; Mège et al., 2021), with an associated uncertainty of approximately 10% (Giannetti et al., 2015).
    If an IRDC lies within an SR, and the SR at the center of the IRDC exhibits the most velocity channels along the line of sight, the SR is classified as near. Due to line-of-sight blending, an IRDC may correspond to multiple SRs at different velocities; however, IRDCs generally correspond to the strongest velocity component (e.g., Kainulainen & Tan, 2013). This approach helps minimize the interference of weaker components by the stronger ones, thereby enhancing the accuracy of the classification.
  • Parameter relations method (7): Distance determination based on the empirical $\sigma_{v}$-$R$ relation (Larson, 1981) resolves near and far kinematic ambiguities by requiring consistency with established scaling laws, a methodology widely implemented in inner Galactic molecular cloud studies (Dame & Thaddeus, 1985; Solomon et al., 1987; Duarte-Cabral et al., 2021). However, as demonstrated in Section 5.3.1 and corroborated by previous investigations (e.g., Heyer et al., 2009; Heyer & Dame, 2015; Miville-Deschênes et al., 2017; Feng et al., 2024), the composite parameter relation $\sigma_{v}$$\Sigma R$ exhibits reduced scatter and demonstrates enhanced robustness across diverse dynamical environments within the Milky Way. Accordingly, we adopt this scaling relation as the premise for developing near- or far-distance discriminants (Miville-Deschênes et al., 2017).
    Building on obtained SR classifications, we propagate distance solutions to individual clumps through the process detailed in Section 4.2, successfully determining distances for over 96% of clumps in the FGQ. We construct a $\sigma_{v}$-$R\Sigma$ distribution plane using samples with known distances as described in Section 5.3.1. For each clump with undetermined classification, we calculate the Mahalanobis distance (Virtanen et al., 2020) between its derived ($\sigma_{v}$, $R\Sigma$) parameters and the multivariate distributions corresponding to both near and far kinematic solutions. This statistical metric, superior to Euclidean alternatives due to its covariance-matrix normalization, effectively accounts for intrinsic parameter correlations in the $\sigma_{v}$-$R\Sigma$ plane and accommodates varying measurement uncertainties. We then adopt the kinematic solution that minimizes this metric.
@@
Figure 8: Face-on number density distribution map of clumps in the Galactic plane. The coordinate system centers on the Galactic center (red star) with the Sun positioned at $(0,8.15)$ kpc (yellow circle). The red dashed-dot circle represents the 3 kpc ring around the Galactic Center, while the blue dashed-dot circle represents the solar circle. The number density distribution is calculated using KDE, where the color scale represents the density of clumps, with higher density regions depicted in greater estimated values. Gray dashed lines show Galactic longitude lines (10° intervals). Axes are centered at the Galactic center, with the negative $y$-axis toward $l=0^{\circ}$ and the positive $x$-axis toward $l=90^{\circ}$. The colored dashed lines depict the spiral arms, with colors matching those in the bottom panel of Figure 2, based on the model by Reid et al. (2019).